ISRAEL JOURNAL OF MATHEMATICS

TABLE OF CONTENTS, VOLUME 132, 2002

Artstein, S. Proportional concentration phenomena on the sphere  337

Beattie, M., S. Dăscălescu and Ş. Raianu Lifting of Nichols algebras of type B_{2}

1

Blanc, D. and G. Peschke The plus construction, Postnikov towers and universal central module extensions

109

Borwein, J. M., M. Fabian and P. D. Loewen The range of the gradient of a Lipschitz C^{1}-smooth bump in infinite dimensions

239

Chan, D., Q. S. Wu and J. J. Zhang Pre-balanced dualizing complexes

285

Clozel, L. Automorphic forms and the distribution of points on odddimensional spheres

175

Dăscălescu, S. See M. Beattie

Defant, A., M. Mastyło and C. Michels Orlicz norm estimates for eigenvalues of matrices

45

Detomi, E., A. Lucchini and F. Morini How many elements are needed to generate a finite group with good probability?

29

Elek, G. Amenable groups, topological entropy and Betti numbers  315
Erschler-Dyubina, A. On the geometry of infinite cyclic subgroups 373

Fabian, M. See J. M. Borwein

Girondo, E. and G. González-Diez Genus two extremal surfaces: extremal discs, isometries and Weierstrass points

221

González-Diez, G. See E. Girondo

Graczyk, P. and C. R. E. Raja Classical theorems of probability on Gelfand pairs — Khinchin’s theorems and Cramér’s theorem

61

Kozlov, D. N. Topology of spaces of hyperbolic polynomials and combinatorics of resonances

189

Lacroix, Y. Possible limit laws for entrance times of an ergodic aperiodic dynamical system

253

Li, G. Ideal Einstein, conformally flat and semi-symmetric immersions 207

Li, Y. On the normalizer problem for integral group rings of torsion groups

169

Loewen, P. D. See J. M. Borwein Lucchini, A. See E. Detomi Mastyło, M. See A. Defant Mattarei, S. The orders of nonsingular derivations of modular Lie

algebras

265

Michels, C. See A. Defant Morini, F. See E. Detomi Mouat, R. and S. Tuncel Constructing finitary isomorphisms with

finite expected coding times

359

Navarro, G. Sylow normalizers and character tables, II  277

Peschke, G. See D. Blanc Pinchover, Y. and T. Saadon (Suez) On positivity of solutions

of degenerate boundary value problems for second-order elliptic equations

125

Raianu, Ş. See M. Beattie Raja, C. R. E. See P. Graczyk Saadon (Suez), T. See Y. Pinchover Tuncel, S. See R. Mouat Wu, Q. S. See D. Chan Zhang, J. J. See D. Chan